OpenStudy (pagen13):

Determine the most efficient way to use the Binomial Theorem to show the following. (11)^4=14641 a.write 11-5+6 and expand. b.write 11=10+1 and expand. c.write 11=3+3+2 and expand. d.write 11=4+4+3 and expand.

OpenStudy (pagen13):

@JoeDeWise

OpenStudy (anonymous):

here

OpenStudy (pagen13):

Okay! I posted my question

OpenStudy (anonymous):

114 =11*11*11*11 =114 =14641

OpenStudy (pagen13):

So?

OpenStudy (freckles):

I would prefer myself to evaluate powers of 10 or powers of 1 because they are extremely easy.

OpenStudy (pagen13):

What do you mean?

OpenStudy (freckles):

for example which one seems easier? 5^4? 10^4? also c doesn't make sense because 3+3+2 is totally not 11 also 11=4+4+3 seems like it might be harder to use binomial theorem since 4+4+3 consist of 3 terms. But I guess you do binomial again and again to expand (4+4+3)^4.

OpenStudy (freckles):

what do I mean about what?

OpenStudy (pagen13):

10^4 seems easier

OpenStudy (freckles):

right

OpenStudy (freckles):

10^4=10000 but 5^4 I would have to go back to multiplying 5*5*5*5 to figure out

OpenStudy (pagen13):

Would it be B?

OpenStudy (freckles):

yes in my opinion expanding (10+1)^4 using binomial theorem is tons easier because like I said before evaluating powers of 10 and evaluating powers of 1 are way easy

OpenStudy (freckles):

\[(a+b)^4=a^4+4 \cdot a^3 b+6 \cdot a^2b^2 + 4 \cdot ab^3+b^4 \\ \text{ so } (10+1)^4=10^4+4 \cdot 10^3(1)+6 \cdot 10^2(1)^2+4 \cdot 10(1)^3+1^4\] try evaluating to see how easy that is

OpenStudy (pagen13):

It was B.! Thank you! Could you help me with three more?

OpenStudy (freckles):

I can try

OpenStudy (pagen13):

Use Pascal's Triangle to expand the binomial (d-3)^6

OpenStudy (freckles):

well using pascal's triangle what are the coefficients in the (6+1)th row?

OpenStudy (pagen13):

I am unsure

OpenStudy (freckles):

so you don't know where the 7th row is?

OpenStudy (pagen13):

I don't understand Pascal's Triangle..

OpenStudy (freckles):

 here I wrote 6 rows